Drops of water fall  from the roof of a building 9m high at regular intervals of time the first drop reach the ground at the same instant fourth drop start its fall. What are the distance of the second and third drop from the roof? 

Grade:12th pass

1 Answers

Khimraj
5 years ago
Let the time interval between falling two drops from the roof be t seconds 
ATQ , 
a drop take 3t seconds from roof to reach the ground 
9 = (1/2)g(3t)² ⇒ t² ⇒ 0.2 
d(2nd) drop = 1/2(g)(2t)² =  4m
d(3rd) drop = 1/2(g)(t)² = 1m
distance = 4-1 =3m
Hope it clears..................